OpenStudy (anonymous):

Find an equation of the line in parametric form of intersection of the two planes x+2y+3z=7 and 2x+x+4z=12.

OpenStudy (anonymous):

So we will have vectors <1,2,3> and <2,1,4>, find the cross product of those first.

OpenStudy (anonymous):

okay, let me do that real quick.

OpenStudy (anonymous):

<5,2,-3>

OpenStudy (anonymous):

Now, set one of the variables in your initial vectors equal to zero and then you will be left with two systems of equations to solve for the remaining two variables.

OpenStudy (anonymous):

x=4+5t y=2t z=1-3t